# Concentration of Measure for Distributions Generated via Diffusion Models

### v2

@@ 4 Experiments
Figure 2 presents some samples from our dataset. See Appendix Figure 20 for more samples and Figure 9 for the spectra of the covariance matrices per class.
Figure 2: Samples from conditional diffusion model with ADM architecture Dhariwal and Nichol [2021] using the checkpoint and sampler settings from Karras et al. [2022], trained on Imagenet64.
Figure 3: The distribution of the $\ell_{2},\ell_{4}$ and $\ell_{10}$ norms of diffusion-generated images for 20 classes of Imagenet64 in our experiments, computed over 10240 samples per class.
4.1 Norm Evolution
We empirically investigate the concentration of norms throughout the sampling process. Following the recommendations of Karras et al. [2022], the diffusion sampling process of $N=256$ steps, with $\sigma(t)=t$, begins with isotropic Gaussian noise of scale $t_{max}:=t_{0}=80$. The noise schedule is constructed as $t_{i<N}=\left({t_{max}}^{\frac{1}{\rho}}+\frac{i}{N-1}\left({t_{min}}^{\frac{1}{\rho}}-{t_{max}}^{\frac{1}{\rho}}\right)\right)^{\rho}$ with $t_{min}:=t_{N-1}=0.002$ and final noise scale $t_{N}=0$. Here, $\rho=7$ is a hyperparameter observed to improve image quality.
Figure 4: The evolution of the $\ell_{2}$ norms through the stochastic sampling process 1. The noise schedule and sampler settings are as prescribed in Karras et al. [2022].
Figure 5: The difference in $\ell_{2}$ norms of intermediate images between consecutive steps of the EDM sampling process. Figure (a) shows the evolution of $||\mathbf{x}^{(i-1)}||-||\mathbf{x}^{(i)}||$ ($i$ denotes sampling step) vs. noise scale for 5000 generation trajectories across 5 classes and (b) shows the mean with variance envelopes of the trajectories, computed over 2048 samples per class. The sampling process clearly appears to be a contraction, supporting Empirical Observation 1
We present further empirical observations regarding the sampling process in Appendix Section D. In Figure 13 we show how the sampling process progressively matches the eigenvalues of the Gram matrix. In Figure 14 we investigate the evolution of the norms of individual pixels.
4.2 Linear Classifier Experiments
We train linear classifiers on our dataset of diffusion-generated images and on the corresponding Gaussian data sampled from a GMM fitted on 10240 diffusion-generated images per class. Following the setting of subsection 3.3, we use SGD as our optimizer and mean squared error (MSE) as our loss criterion. For multi-class classification, we use a softmax activation on the logits and compute the MSE loss against the one-hot-encoded class labels. For binary classification, we compute the MSE loss on the logit after sigmoid activation. This regression on predicted class probabilities is done in practice when working with soft (noisy) labels or in the context of knowledge distillation.
We compare the accuracies achieved by linear classification on the diffusion-generated images versus the GMM samples, when varying the size of the training set between 128 and 4096 samples per class. Figures 6 and 7 present the results of 10-20 independent runs per training data split, with a different random seed for each run. Thus, for each run, a unique pseudorandom generator state determines weight initialization in addition to the sampling and minibatch shuffling of $N_{\text{train per class}}\in[128,4096]$ samples from our dataset of 10240 samples per class for diffusion-generated images and GMM samples. Likewise, we fix the size of our held-out test set to $N_{\text{test per class}}=1024$, randomly sampled according to each run’s unique random state from a separate subset of our dataset to ensure no overlap with the training set.
Figure 6: Linear classifier accuracies for diffusion-generated images (Red) and GMM samples (Blue). We consider 4, 10, and 20 class mixtures of Baseball, Cauliflower, Church, Coral Reef, English Springer, French Horn, Garbage Truck, Goldfinch, Kimono, Mountain Bike, Patas Monkey, Pizza, Planetarium, Polaroid, Racer, Salamandra, Tabby, Tench, Trimaran, Volcano.
Figure 7: Binary linear classifier accuracies for diffusion-generated images (Red) and GMM samples (Blue).
To robustly achieve the best possible linear classification performance, we perform an extensive sweep over a range of batch sizes and of learning rates between $[10^{-4},0.1]$, while ensuring convergence with respect to the test loss, and no overfitting. For practical reasons, we use a cosine annealing learning rate schedule to speed up convergence.
In summary, we observe matching accuracies for linear classifiers trained on diffusion-generated images and on the corresponding Gaussian data, for a range of training set sizes and class subsets. At large values of $N_{\text{train per class}}$ we begin to see a divergence in the accuracies, which we attribute to the estimation gap from computing the mean and covariance from a finite number of samples. The choice of MSE loss on one-hot-encoded labels may seem unconventional for classification but is done to match the setting of 3.3. We re-ran our experiments using cross-entropy loss and also observe a match, but did not conduct as extensive a sweep for this setting and expect the match to improve.
Figure 8: Accuracies for linear classifier trained with cross-entropy loss, for diffusion images (Red) and GMM (Blue).
In Appendix Section C we compute the eigenvalue spectra of the Gram matrices of multi-class mixtures of diffusion-generated images. Figures 11 and 12 in the appendix show a very close match between the Gram spectra of diffusion-generated images and that of the corresponding GMM. And in Appendix Section E we investigate the Gram spectra and eigenspaces of ResNet representations of high-resolution images sampled from a latent-diffusion model, showing a close match (Figures 15, 16, and 17) and suggesting an investigation into the concentration of latent diffusion models and their representations as an avenue for future work.
5 Conclusion
In this work, we focus on characterizing the mathematical properties of images generated by diffusion models by examining the generalization error of the classification tasks. We are motivated by the fact that characterizing the generalization error and performance of neural networks precisely remains one of most challenging problems in modern machine learning. In fact, most theoretical works have focused on analyzing models under specific assumptions about data distribution, such as isotropic Gaussianity even though real-world datasets are almost never Gaussian. As such, we choose to study theoretical properties of the diffusion-generated distributions instead as an approximation to real-world distributions more amenable to analyses. Future directions include extending the universality results to accommodate for more general covariance matrices, incorporating training with softmax into the universality framework and providing a rigorous proof of the contractivity of the sampling process.
Figure 9: Spectra of covariance matrices of diffusion-generated images. Scaled by exponent $0.1$ for presentation.
Figure 10: First 1000 eigenvalues of the covariance matrices.
6 Acknowledgments
We would like to thank Joel A. Tropp for insightful discussions and Morteza Mardani for suggesting to look into the evolution of the norms of the individual pixels during the sampling process. AB is grateful for the support of the Basdall Gardner Memorial MCD Fellowship and for the gracious maintenance of computational resources carried out by the Research Computing Task Force (RCTF) at UT Austin.
Appendix A Proof of Theorem 1
First note that we will utilize a multi-dimensional version of the CLT result of Bobkov [2003] (Corollary 2.5) which controls the following quantity for a matrix $\mathbf{W}$ with "generic" column vectors:
$\biggl{|}\mathbb{P}\Bigl{(}i\neq\operatorname*{arg\,max}_{\ell\in[k]}\mathbf{W}_{\ell,\Phi_{\lambda}(\mathbf{X})}^{T}\mathbf{x}\Bigl{|}\mathbf{x}\sim\mathbb{P}_{i}\Bigr{)}-\mathbb{P}\Bigl{(}i\neq\operatorname*{arg\,max}_{\ell\in[k]}\mathbf{W}_{\ell,\Phi_{\lambda}(\mathbf{X})}^{T}\mathbf{g}\Bigl{|}\mathbf{x}\sim\mathbb{P}_{i}\Bigr{)}\biggr{|}$
This generalization follows by applying a union bound argument. For the main quanitity of interest,
$\biggl{|}\mathbb{P}\Bigl{(}i\neq\operatorname*{arg\,max}_{\ell\in[k]}\mathbf{W}_{\ell,\Phi_{\lambda}(\mathbf{X})}^{T}\mathbf{x}\Bigl{|}\mathbf{x}\sim\mathbb{P}_{i}\Bigr{)}-\mathbb{P}\Bigl{(}i\neq\operatorname*{arg\,max}_{\ell\in[k]}\mathbf{W}_{\ell,\Phi_{\lambda}(\mathbf{G})}^{T}\mathbf{g}\Bigl{|}\mathbf{x}\sim\mathbb{P}_{i}\Bigr{)}\biggr{|}$
We would need to bound the following:
$\biggl{|}\mathbb{P}\Bigl{(}i\neq\operatorname*{arg\,max}_{\ell\in[k]}\mathbf{W}_{\ell,\Phi_{\lambda}(\mathbf{X})}^{T}\mathbf{g}\Bigl{|}\mathbf{x}\sim\mathbb{P}_{i}\Bigr{)}-\mathbb{P}\Bigl{(}i\neq\operatorname*{arg\,max}_{\ell\in[k]}\mathbf{W}_{\ell,\Phi_{\lambda}(\mathbf{G})}^{T}\mathbf{g}\Bigl{|}\mathbf{x}\sim\mathbb{P}_{i}\Bigr{)}\biggr{|}$
Which involves analyzing the covariance and the mean of $\mathbf{W}_{\ell,\Phi_{\lambda}(\mathbf{A})}^{T}\mathbf{g}$ for $\mathbf{A}=\mathbf{G},\mathbf{X}$.
We know from Thrampoulidis et al. [2020] for the case of a GMM, (see Equation 2.7 in Thrampoulidis et al. [2020]) that the generalization error is characterized by the quantities $\bm{\mu}_{\ell}^{T}(\mathbf{w}_{\ell}-\mathbf{w}_{\ell^{\prime}})$, and $\bm{\Sigma}^{1/2}_{\ell}\mathbf{S}\bm{\Sigma}^{1/2}_{\ell}$ where $(S_{\ell})_{ij}:=(\mathbf{w}_{i}-\mathbf{w}_{j})^{T}(\mathbf{w}_{i}-\mathbf{w}_{j})$ for $i,j\neq\ell$. Consider the following ridge regression objective:
$\Phi_{\lambda}(\mathbf{A}):=\min_{\mathbf{W}}\frac{\lambda}{2}\|\mathbf{A}\mathbf{W}-\mathbf{Y}\|_{F}^{2}+\|\mathbf{W}\|_{F}^{2}=\sum_{\ell=1}^{k}\min_{\mathbf{w}_{\ell}}\frac{\lambda}{2}\|\mathbf{A}\mathbf{w}_{\ell}-\mathbf{y}_{\ell}\|_{2}^{2}+\|\mathbf{w}_{\ell}\|_{2}^{2}$
We denote the solution to the above optimization problem as $\mathbf{W}_{\Phi_{\lambda}(\mathbf{A})}$. Now in order to characterize $(S_{\ell})_{ij}$, note that we need to understand the pairwise interaction of $\mathbf{w}_{i}$ and $\mathbf{w}_{j}$ and the decomposition provided cannot capture these quantities. To do this, we use the following identity:
$=\min_{\mathbf{w}_{i}-\mathbf{w}_{j},\mathbf{w}_{i}+\mathbf{w}_{j}}\frac{\lambda}{4}\|\mathbf{A}(\mathbf{w}_{i}+\mathbf{w}_{j})-\mathbf{y}_{\ell}\|_{2}^{2}+\|\mathbf{w}_{i}+\mathbf{w}_{j}\|_{2}^{2}+\frac{\lambda}{4}\|\mathbf{A}(\mathbf{w}_{i}-\mathbf{w}_{j})-\mathbf{y}_{\ell}\|_{2}^{2}+\|\mathbf{w}_{i}-\mathbf{w}_{j}\|_{2}^{2}$
And by studying the norms of $\bm{\Sigma}_{\ell}^{1/2}(\mathbf{w}_{i}\pm\mathbf{w}_{j})$ we can understand $(S_{\ell})_{ij}$. Note that in the argument above $\mathbf{A}$ could be either $\mathbf{X}$ and $\mathbf{G}$ as a multi-dimensional CLT argument reduces the problem of universality of the test error on $\mathbf{X}$ to $\mathbf{G}$ and it only requires the first and second order statistics of $\mathbf{X}$. Note that by combining the results of Ghane et al. [2024] and the above identity, we observe that $|(S_{\ell}(\mathbf{X}))_{ij}-(S_{\ell}(\mathbf{G}))_{ij}|\xrightarrow[]{\mathbb{P}}0$ for every $\lambda$. We observe that if $\Phi_{\lambda}(\mathbf{A})\xrightarrow[]{\mathbb{P}}c_{\lambda}$ then $\sup_{\lambda>0}\Phi_{\lambda}(\mathbf{A})\xrightarrow[]{\mathbb{P}}\sup_{\lambda>0}c_{\lambda}$. Combining this with a perturbation argument concludes the proof.
Appendix B Proof of Theorem 2
We rely on the following proposition, claiming that the image of an exponentially concentrated distribution under a Lipschitz transformation remains exponentially concentrated, albeit with different constants. It follows trivially from the definition of concentration but is nonetheless crucial for our purposes:
Proposition 2.
Let $\mathbf{x}$ satisfy Definition 1 with constants $C,\sigma$ and assume that $\mathcal{G}:\mathbb{R}^{d}\to\mathbb{R}^{D}$ is $L$- Lipschitz. Then $\mathcal{G}(\mathbf{x})$ satisfies CoM from Definition 1 as well with constants $C,L\sigma$.
Now, recall that the images are generated step by step according to:
$\mathbf{x}_{0}\approx\mathbf{x}^{(N)}=\mathcal{R}^{(N-1)}_{D_{\theta}}(\mathcal{R}^{(N-2)}_{D_{\theta}}(\ldots\mathcal{R}^{(0)}_{D_{\theta}}(\mathbf{x}^{(0)},t_{0:1})\ldots),t_{N-2:N-1}),t_{N-1:N})$ (6)
We will prove that each $\mathbf{x}_{i}$ satisfies the CoM of measure property by induction by $i=0,\dots,N$:
$\bullet$ Basis $i=0$ follows from Proposition 1
$\bullet$ Step $i\to i+1$ would follow by applying Proposition 2 with $L=1$ to $\mathbf{x}^{(i+1)}=\mathcal{R}^{(i)}_{D_{\theta}}(\mathbf{x}^{(i)},t_{i:i+1})$ if we knew that $\mathcal{R}^{(i)}_{D_{\theta}}(.,t_{i:i+1})$ is $1$-Lipschitz. Since we already assume that $\mathcal{R}^{(i)}_{D_{\theta}}$ is norm-decreasing, it suffices to just prove that it is Lipschitz. The latter does not have to hold in general but in this case we make a specific assumption that $\mathcal{R}^{(i)}_{D_{\theta}}$ corresponds to a denoiser employing a U-Net architecture. This is a neural network consisting of the following blocks:
@@ Appendix B Proof of Theorem 2
  • Self-Attention Layers As shown in Kim et al. [2021], these are not Lipschitz over the entire domain. However, it can be seen from the same derivations that, if we restrict the domain to points from a distribution satisfying CoM, then it is Lipschitz with high probability. Thus, for our specific scenario, these are Lipschitz as well.
@@
We conclude that the mapping $\mathbf{x}^{(i)}\to\mathbf{x}^{(i+1)}$ is Lipschitz but, technically speaking, the constant is unbounded without the assumption that $\mathcal{R}^{(i)}_{D_{\theta}}$ does not increase norms. Moreover, since the sampling process involves $N$ steps for a relatively big $N$, the Lipschitz constant of the mapping $\mathbf{x}_{T}\to\mathbf{x}_{0}$ might accumulate and explode unless the Lipschitz constant of each step is bounded by $1$. While we could not prove directly that the latter is the case so far, we observed it to be the case in the simulations we have conducted (cf. Figure 4). As such, we decided to assume that the training is performed in such a manner that the sampling steps $\mathbf{x}^{(i)}\to\mathbf{x}^{(i+1)}$ are all $1$-Lipschitz mappings for the scope of the present work. In addition, we can prove that the same conclusion holds if we only assume that each $\mathcal{R}_{i}$ is norm-decreasing for large enough $t>K$ for a constant $K$. This follows automatically from the following result, that we could not find in the literature:
Theorem 3.
If $(x,y)\sim\Pi$ where $\Pi$ is a joint distribution with marginals, $\pi_{1\#}\Pi=p_{1}$ and $\pi_{2\#}\Pi=p_{2}$ and $p_{1}$ and $p_{2}$ are two distributions satisfying CoM, then $(x,y)$ also satisfies CoM.
@@
The proof technique is adopted from Ledoux [2006]. We also extend that result to accommodate for arbitrary coupling on $(x,y)$. For every $L$-Lipschitz function $f:\mathbb{R}^{d}\times\mathbb{R}^{d}\rightarrow\mathbb{R}$, we have from triangle inequality
$\mathbb{P}(|f(x,y)-\mathbb{E}_{\Pi}f(x,y)|>2t)\leq\mathbb{P}(|f(x,y)-\mathbb{E}_{p_{1}}f(x,y)|>t)+\mathbb{P}(|\mathbb{E}_{p_{1}}f(x,y)-\mathbb{E}_{\Pi}f(x,y)|>t)$ (7)
For the first term in 7, we have that for the joint distribution $\Pi$
$\mathbb{P}(|f(x,y)-\mathbb{E}_{p_{1}}f(x,y)|>t)$ $=\mathbb{E}_{\Pi}\mathds{1}\{|f(x,y)-\mathbb{E}_{p_{1}}f(x,y)|>t\}=\mathbb{E}_{p_{2}}\mathbb{E}_{p_{1}|p_{2}}\mathds{1}\{|f(x,y)-\mathbb{E}_{p_{1}}f(x,y)|>t\}$
$=\mathbb{E}_{p_{2}}\mathbb{P}_{p_{1}|p_{2}}(|f(x,y)-\mathbb{E}_{p_{1}}f(x,y)|>t)$
Now we observe that for every $y$, $f(x,y)$ is also $L$-Lipschitz in $x$, thus by CoM
$\mathbb{P}(|f(x,y)-\mathbb{E}_{p_{1}}f(x,y)|>t)=\mathbb{E}_{p_{2}}\mathbb{P}_{p_{1}|p_{2}}(|f(x,y)-\mathbb{E}_{p_{1}}f(x,y)|>t)\leq\mathbb{E}_{p_{2}}Ce^{-(\frac{t}{L\sigma})^{2}}=Ce^{-(\frac{t}{L\sigma})^{2}}$
For the second term in 7, letting $g(y):=\mathbb{E}_{p_{1}}f(x,y)$, we observe that $g$ is also Lipschitz, so by CoM for $p_{2}$,
$\mathbb{P}(|\mathbb{E}_{p_{1}}f(x,y)-\mathbb{E}_{\Pi}f(x,y)|>t)\leq Ce^{-(\frac{t}{L\sigma})^{2}}$
@@
$\mathbb{P}(|f(x,y)-\mathbb{E}_{\Pi}f(x,y)|>2t)\leq 2Ce^{-(\frac{t}{L\sigma})^{2}}$
Appendix C Gram Spectrum
For each of the subsets of classes we considered for our multiclass linear classification experiments, we also investigated the spectrum of the gram matrix of the corresponding mixture distribution. Using an equal number of samples per class, we construct a data matrix $X\in\mathbb{R}^{n\times d}$ where $n$ is the total number of samples and $d=12288$ is the dimensionality of each sample (viewed as a vector). Figure 11 presents the eigenvalue spectrum of the resulting Gram matrix of the type $XX^{T}\in\mathbb{R}^{n\times n}$. As can be seen, we observe a very close match between the distributions of the eigenvalues of the Gram matrices for the diffusion-generated data and the corresponding GMM, but there is a slight mismatch for the smaller eigenvalues. We leave the question of finding out if there are any reasons for the latter mismatch apart from numerical inaccuracies for future work.
Figure 11: Spectra of Gram Matrices for balanced mixtures of samples from 4, 10, and 20 classes considered in the linear classification experiments. Computed for Diffusion (Red) and GMM (Blue). We use 2048 samples per class for 4 classes, and 512 samples per class for 10 and 20 classes.

### v3

@@ Appendix D Application of Theorem 2 to conventional samplers
We also test the validity of our results empirically for the U-Net architecture because the latter is often employed in practice. Despite our prediction matches the empirical results well for this architecture too, note that we were not able to verify the high-probability Lipschitness assumptions for this network, leaving it as an interesting challenge that will hopefully stimulate future work in this direction. To be precise, U-Net is a neural network consisting of the following blocks:
@@
  • Self-Attention Layers As shown in Kim et al. [2021], these are not Lipschitz over the entire domain. However, intuitively it should be true that, if we restrict the domain to points sampled from a distribution satisfying ACoM, then it is Lipschitz with high probability. We have not been able to verify it by the moment of submission and encourage an interested reader to do so. To support our point further, if this architecture were not Lipschitz for the inputs of interest with high probability, one would imagine that it would be utterly impractical to train and use. However, nobody has reported this for self-attention.
@@
We conclude that the mapping $\mathbf{x}^{(i)}\to\mathbf{x}^{(i+1)}$ is Lipschitz but, technically speaking, the constant is unbounded. Moreover, since the sampling process involves $N$ steps for a relatively large $N$, the Lipschitz constant of the mapping $\mathbf{x}_{T}\to\mathbf{x}_{0}$ might accumulate and explode unless the Lipschitz constant of each step is bounded by $1$. While we could not prove directly that the latter is the case so far, we observed $\mathcal{R}^{(i)}_{D_{\theta}}$ to be contractive in the simulations we have conducted (cf. Figure 3). As such, we decided to assume that the training is performed in such a manner that the sampling steps $\mathbf{x}^{(i)}\to\mathbf{x}^{(i+1)}$ are all $1$-Lipschitz mappings for the scope of the present work.
Lemma 2.
If $(x,y)\sim\Pi$ where $\Pi$ is a product measure with marginals, $\pi_{1\#}\Pi=p_{1}$ and $\pi_{2\#}\Pi=p_{2}$ , $p_{1}$ and $p_{2}$ are two distributions satisfying ACoM with $\Bigl(C_{1},c_{1},c^{\prime}_{1},d,\sigma_{1}\Bigr)$ and $\Bigl(C_{2},c_{2},c^{\prime}_{2},d,\sigma_{2}\Bigr)$, respectively, then $(x,y)$ also satisfies ACoM with $\Bigl(C_{1}+C_{2},c_{1}+c_{2},\min\{c^{\prime}_{1},c^{\prime}_{2}\},d,\max\{\sigma_{1},\sigma_{2}\}\Bigr)$.
@@
The proof technique is adopted from Ledoux [2001]. For every $L$-Lipschitz function $f:\mathbb{R}^{d}\times\mathbb{R}^{d}\rightarrow\mathbb{R}$, we have from triangle inequality
$\mathbb{P}\Bigl(|f(x,y)-\mathbb{E}_{\Pi}$ $f(x,y)|>2t\Bigr)$
$\leq\mathbb{P}\Bigl(|f(x,y)-\mathbb{E}_{p_{1}}f(x,y)|>t\Bigr)$
$+\mathbb{P}\Bigl(|\mathbb{E}_{p_{1}}f(x,y)-\mathbb{E}_{\Pi}f(x,y)|>t\Bigr)$ (6)
For the first term in D, we have that for the product measure $\Pi$
$\mathbb{P}\Bigl(|f(x,y)$ $-\mathbb{E}_{p_{1}}f(x,y)|>t\Bigr)$
$=\mathbb{E}_{\Pi}\mathds{1}\Bigl\{|f(x,y)-\mathbb{E}_{p_{1}}f(x,y)|>t\Bigr\}$
$=\mathbb{E}_{p_{2}}\mathbb{E}_{p_{1}|p_{2}}\mathds{1}\Bigl\{|f(x,y)-\mathbb{E}_{p_{1}}f(x,y)|>t\Bigr\}$
$=\mathbb{E}_{p_{2}}\mathbb{P}_{p_{1}|p_{2}}\Bigl(|f(x,y)-\mathbb{E}_{p_{1}}f(x,y)|>t\Bigr)$
Now we observe that for every $y$, $f(x,y)$ is also $L$-Lipschitz in $x$, thus by ACoM
$\mathbb{P}\Bigl(|f(x,y)-\mathbb{E}_{p_{1}}$ $f(x,y)|>t\Bigr)$
$=\mathbb{E}_{p_{2}}\mathbb{P}_{p_{1}|p_{2}}\Bigl(|f(x,y)-\mathbb{E}_{p_{1}}f(x,y)|>t\Bigr)$
$\leq C_{1}e^{-(\frac{t}{L\sigma_{1}})^{2}}+c_{1}e^{-c_{1}^{\prime}d}$
For the second term in D, letting $g(y):=\mathbb{E}_{p_{1}}f(x,y)$, we observe that $g$ is also Lipschitz, so by ACoM for $p_{2}$,
$\mathbb{P}\Bigl(|\mathbb{E}_{p_{1}}f(x,y)-\mathbb{E}_{\Pi}f(x,y)|>t\Bigr)\leq C_{2}e^{-(\frac{t}{L\sigma_{2}})^{2}}+c_{2}e^{-c_{2}^{\prime}d}$
@@
$\mathbb{P}\Bigl(|f(x,y)$ $-\mathbb{E}_{\Pi}f(x,y)|>2t\Bigr)$
$\leq C_{1}e^{-(\frac{t}{L\sigma_{1}})^{2}}+c_{1}e^{-c_{1}^{\prime}d}+C_{2}e^{-(\frac{t}{L\sigma_{2}})^{2}}+c_{2}e^{-c_{2}^{\prime}d}$
$\leq(C_{1}+C_{2})\exp\Bigl(-\Bigl(\frac{t}{L\max\{\sigma_{1},\sigma_{2}\}}\Bigr)^{2}\Bigr)$
$+(c_{1}+c_{2})\exp\Bigl(-\min\{c_{1}^{\prime},c_{2}^{\prime}\}d\Bigr)$
Which concludes the proof. ∎
Appendix E Gram Spectrum
For each of the subsets of classes we considered for our multiclass experiments, we also investigated the spectrum of the Gram matrix of the corresponding mixture distribution. Using an equal number of samples per class, we construct a data matrix $\mathbf{X}\in\mathbb{R}^{n\times d}$ where $n$ is the total number of samples and $d=12288$ is the dimensionality of each sample (viewed as a vector). Figure 6 presents the eigenvalue spectrum of the resulting Gram matrix of the type $\mathbf{X}\mathbf{X}^{T}\in\mathbb{R}^{n\times n}$. As can be seen, we observe a very close match between the distributions of the eigenvalues of the Gram matrices for the diffusion-generated data and the corresponding GMM, but there is a slight mismatch for the smaller eigenvalues. We leave the question of finding out if there are any reasons for the latter mismatch apart from numerical inaccuracies for future work.
Figure 6: Spectra of Gram Matrices for balanced mixtures of 4, 10, and 20 classes, for Diffusion (Red) and GMM (Blue). We use 2048 samples per class for the 4-class mixture, and 512 samples per class for 10 and 20 class mixtures.